ABC Company employs a periodic inventory system and sells its inventory to
customers for $23 per unit. ABC Company reported the following inventory
information for the month of May:

May 1    Beginning inventory 4,200 units @ $9 cost per unit
May 6    Purchased 2,000 units @ $12 cost per unit
May 8    Sold 3,000 units
May 13   Purchased 4,000 units @ $8 cost per unit
May 18   Sold 2,100 units
May 21   Purchased 2,500 units @ $6 cost per unit
May 28   Sold 2,200 units
May 30   Purchased 2,300 units @ $17 cost per unit

ABC Company reported operating expenses of $26,100 for May and they had a
tax rate of 43%.

Calculate the dollar amount of ending inventory shown on ABC Company's
May 31 balance sheet using the FIFO method.
Answer #1
Unit Unit Cost Total cost
Beginning inventory 4200 9 37800
Purchase 1 2000 12 24000
Purchase 2 4000 8 32000
Purchase 3 2500 6 15000
Purchase 4 2300 17 39100
Total 15000 147900

Sales unit = 3000+2100+2200 =7300

Ending inventory unit = 15000-7300 = 7700 Units

Ending inventory = 39100+15000+2900*8 = $77300